m=3/4, (7,-1)
the equation of the line is y=

Answer by texttutoring(324)

Use y=mx+b, with x=7, y=-1 and m=3/4. Solve for b.

y=mx+b
-1 = (3/4)7 + b
-1 = 21/4 + b
-1 - 21/4 = b
-4/4 - 21/4 = b
-25/4 = b

The equation of the line is:

y=(3/4)x -25/4